OpenStudy (anonymous):

A truck undergoes a constant acceleration of 2.0 meters/second2 on a highway that slopes upward at an angle of 30.0° to the horizontal. What is the magnitude of the horizontal component of the acceleration?

OpenStudy (anonymous):

@UnkleRhaukus

OpenStudy (anonymous):

@chmvijay

OpenStudy (unklerhaukus):



OpenStudy (unklerhaukus):

\[\boxed{\cos\theta=\frac{\text{adjacent side}}{\text{hypotenuse}}},\]